# One Model for All Tasks: Leveraging Efficient World Models in Multi-Task Planning

### v1

@@ 3.1 Diagnosing Plasticity Collapse in Multi-Task Learning
Figure 1: Diagnosing plasticity collapse in the baseline UniZero model on a multi-task Atari benchmark. While simple tasks like Pong and Hero show stable learning, complex tasks such as Seaquest and ChopperCommand suffer a catastrophic performance collapse in later training (Top). This performance failure is precisely correlated with a sharp spike in the dormant neuron ratio of the Transformer backbone (Bottom Left) and an uncontrolled inflation of the latent state norm (Bottom Right), empirically validating the link between external performance and internal network dynamics.
Metrics. In MTRL settings, shared-parameter models are susceptible to a freezing phenomenon, where parts of the network become saturated and lose their ability to adapt to new information. This plasticity collapse stems from the inherent heterogeneity across tasks in terms of their features, difficulty, and informational content, ultimately constraining the model’s overall learning potential. To quantitatively measure this degradation in network adaptability, we utilize two key metrics: (1) Dormant Neuron Ratio (Sokar et al., 2023): This metric measures the proportion of neurons in a given layer whose activation magnitude falls below a predefined threshold $\epsilon$ (set to 0.025 in this work) over a time series. A high dormant ratio indicates that a significant portion of the network’s pathways are inactive and no longer contribute to information processing. $\text{DormantRatio}(l)=\frac{1}{N_{l}}\sum_{i=1}^{N_{l}}\mathbf{1}\!\left(|h_{i}^{l}|\leq\epsilon\right),$ where $h_{i}^{l}$ is the activation of the $i$-th neuron in layer $l$, and $N_{l}$ is the total number of neurons in that layer. (2) Latent State Norm: This metric computes the average L2 norm of the latent state vectors $z_{t}$ produced by the encoder over a time series. An abnormally inflated latent norm is often correlated with training instability and slower learning. As our model processes sequential data in mini-batches, we average this norm across both the training sequence length $T$ and the mini-batch size $B$.
Analysis. Our experiments on a multi-task suite of 8 Atari games (see Appendix A) reveal a clear failure pattern: while simple tasks (e.g., Pong) converge stably, complex tasks (e.g., Seaquest) suffer a catastrophic performance collapse late in training (Figure 1). This collapse precisely corresponds to a sharp increase in the dormant neuron ratio and an uncontrolled inflation of the latent state norm as shown in bottom of Figure 1. In addition, we have added the trend of the feature effective rank Dohare et al. (2024) and training return in Appendix Figure 6, which also shows a marked decline in the performance degradation range. Thus, we attribute this plasticity loss to two intertwined factors: (1) Gradient Competition, where simpler, faster-converging tasks generate dominant gradients that suppress signals from more complex tasks, causing critical neural pathways to "go dormant"; and (2) Representation Interference, where a single, shared model struggles to encode diverse task dynamics, creating a representational bottleneck that constricts the solution space.
Motivated by this diagnosis, we propose to mitigate these issues by introducing task-differentiated mechanisms from two distinct perspectives. First, from the perspective of a single learning iteration, we investigate modifications that resolve interference within each forward and backward pass, leading to the design of ScaleZero (Section 3.2). Second, from the perspective of the overall learning process, we introduce a novel training strategy that dynamically allocates model capacity over time, we named as Dynamic Parameter Scaling strategy (Section 3.3).
3.2 Exploring the Design Space of UniZero
@@
Figure 2: (a) A systematic exploration of the UniZero design space across five axes: task conditioning, encoder architecture, latent normalization, backbone design, and multitask optimization. This investigation informs the design of our proposed ScaleZero model. (b) A conceptual diagram of Dynamic Parameter Scaling (DPS), our adaptive training strategy. DPS progressively expands model capacity by injecting lightweight LoRA adapters in stages, triggered by learning progress, to direct computational resources toward unsolved tasks while preserving existing knowledge.
@@
Figure 3: Performance impact of architectural modifications on the Atari8 multi-task benchmark. This ablation study across the UniZero design space reveals that replacing the standard Transformer backbone with a Mixture-of-Experts architecture yields the most significant and consistent performance gains. In contrast, other interventions, with the partial exception of SimNorm, provide marginal or inconsistent benefits. These results underscore the centrality of the MoE’s conditional computation in overcoming the limitations of a shared, dense backbone.
Figure 4: Detailed analysis correlating performance with plasticity metrics across model variants in Figure 3. On two representative hard tasks (ChopperCommand and Seaquest), the performance collapse of baseline models directly correlates with catastrophic plasticity loss—evidenced by a sharp rise in the dormant neuron ratio and latent norm. In stark contrast, the MoE-based model sustains high returns by keeping the dormant ratio near-zero and the latent norm stable. This confirms that the MoE’s superior performance stems directly from its success in mitigating the diagnosed plasticity issues. Analyses for the remaining six environments, which show consistent trends, are deferred in Figure 8.
Based on the diagnosis above, we conducted a systematic decomposition and evaluation of the UniZero model’s design space to identify architectural components that can effectively counteract plasticity loss. We explored five key dimensions, as illustrated in Figure 2 (left): task information, encoder, latent normalization, backbone, and optimization strategy. The experimental details is showed in Appendix B. Figure 3 provides a overall comparison of the performance contributions from each design dimension. A key observation is that the MoE backbone yields the most significant performance gains. In contrast, other modifications, with the partial exception of SimNorm, fail to deliver stable performance improvements. To investigate the mechanisms behind these results, Figure 4 illustrates the relationship between task returns and plasticity-related metrics (i.e., transformer dormant ratio and latent norm) for two representative tasks, ChopperCommand and Seaquest. Analyses for the remaining six environments are provided in Appendix B. The key findings from this exploration are detailed below:
Explicit Task Conditioning. We experimented with concatenating a learnable task embedding indexed by task ID to the state representation to provide an explicit task identity. While this accelerated initial convergence on simpler tasks, its long-term benefit for complex tasks was marginal, and it failed to significantly improve internal plasticity metrics (Figure 4). This suggests that shallow task identifiers are insufficient to resolve deep-seated representational conflicts in online RL setting.
Encoder Architecture (ResNet vs. ViT). We compared a standard ResNet encoder against a ViT to assess the impact of encoder architecture. In our experiments on the Atari8 benchmark, the choice of encoder did not yield a significant performance advantage. We hypothesize that this is due to the limited number of tasks and the relatively low visual heterogeneity across the environments in this specific benchmark. However, we anticipate that the ViT’s superior capacity and scalability will become more critical in larger-scale, more heterogeneous multi-task settings.
Latent Normalization (LayerNorm vs. SimNorm). To suppress the inflation of the latent norm, we compared a standard LayerNorm (Ba et al., 2016) against a SimNorm (Hansen et al., 2023) to assess the impact of latent normalization. SimNorm (Hansen et al., 2023) normalize the latent representation by projecting $z$ into $L$ fixed-dimensional simplices using a softmax operation, which clamps the L2 norm of latent vectors within a predefined threshold. Although SimNorm effectively stabilized training and prevented performance collapse, its hard constraint appeared to curtail the expressive power of the representations, highlighting a trade-off between stability and representational capacity.
Mixture-of-Experts Backbone. To fundamentally address plasticity loss in the Transformer backbone, we replaced its standard feed-forward networks (FFNs) with MoE layers. In an MoE, a gating network sparsely routes each input token to a subset of "expert" sub-networks. This conditional computation mechanism decouples the dense computational pathways into multiple, parallel, specialized ones. This design: (1) reduces parameter collision, as different tasks can leverage different combinations of experts, mitigating representation interference; and (2) mitigates gradient conflict, as gradient updates for different tasks are largely directed to the distinct sets of activated experts. As shown in Figures 4 and 4, the MoE backbone delivered the most significant performance improvements on complex tasks like Seaquest and maintained a relatively low dormant ratio throughout training, effectively controlling the latent norm.
@@
Multi-Task Gradient Correction. We implemented a dynamic gradient re-weighting scheme inspired by MoCo (Fernando et al., 2023) to directly address gradient interference. Despite yielding gains on some tasks, its significant computational overhead and inconsistent efficacy across the heterogeneous task suite led us to explore a more efficient gradient correction in future.
ScaleZero. Based on this exploration, we integrated the most impactful components to form the ScaleZero architecture. Its core design combines: (1) a ViT Encoder for powerful feature extraction; (2) an MoE Backbone to fundamentally address plasticity collapse through sparse, conditional computation; and (3) Standard Layer Normalization on the latent states to ensure stability without sacrificing expressive power. As validated in Section 4, ScaleZero provides a novel and effective blueprint for building generalist world models.
3.3 Dynamic Parameter Scaling for Balanced Multi-Task Learning
In heterogeneous MTRL, static, one-size-fits-all model architectures are suboptimal. They tend to waste computational resources on tasks that have already been mastered and often lack the necessary plasticity to effectively tackle the remaining, more difficult tasks. This can lead to inefficient training and risks negative transfer. To address this, we propose Dynamic Parameter Scaling (DPS), a strategy illustrated in Figure 2 (right). The core insight of DPS is that a model’s effective capacity should dynamically align with the current set of unsolved tasks. This approach instantiates a "curriculum of model complexity," adaptively expanding the model with new parametric resources directed precisely at the most challenging tasks still in progress.
Formalism and Mechanism. Let the set of tasks be $\mathcal{T}=\{\tau_{i}\}_{i=1}^{N}$ and the parameters of the shared backbone network be $\theta_{B}$. We define a task $\tau_{i}$ as "solved" when its performance metric, $\text{Metric}(\tau_{i})$, meets or exceeds a predefined threshold $\varepsilon_{i}$. At any time $t$, the set of all unsolved tasks is denoted as the active set $\mathcal{U}_{t}\subseteq\mathcal{T}$. To conserve computational resources, both data collection and model training are exclusively performed on the tasks within this active set. Once a task is considered solved, it is removed from $\mathcal{U}_{t}$, and we cease all related training activities for it. The training protocol is divided into $S+1$ stages ($s=0,1,\dots,S$). It begins with Stage 0 (Warm-up), an initial period of $T_{0}$ iterations where only the backbone parameters $\theta_{B}$ are trained. Each subsequent Stage $s\geq 1$ (Expansion) introduces an independent Low-Rank Adaptation (LoRA) module, $\Delta\theta_{s}$. This module consists of a trainable matrix pair $\Delta\theta_{s}=B_{s}A_{s}$ (where $B_{s}\in\mathbb{R}^{d\times r},A_{s}\in\mathbb{R}^{r\times k}$ with rank $r\ll d,k$) and a learnable scalar scaling factor $\alpha_{s}\in\mathbb{R}$.
@@ 3.3 Dynamic Parameter Scaling for Balanced Multi-Task Learning
Adaptive Stage Transition Triggers. The transition from stage s1s-1 to stage $s$ is governed by one of two conditions, balancing learning progress against a fixed training budget: (1) Progress-based Trigger: A new stage is initiated when the number of newly solved tasks since the beginning of the current stage reaches a predefined quota $Q_{s}$. (2) Budget-based Trigger: To prevent training from stalling on exceptionally difficult tasks, a budget-based trigger forces a transition if the number of iterations in the current stage exceeds a pre-allocated limit, $\lceil(T_{\max}-T_{0})/S\rceil$.
Parameter Composition and Optimization. In any given stage $s$, the effective computation for a weight matrix $W_{0}\in\theta_{B}$ is expressed as $W^{(s)}(x)=W_{0}x+\sum_{j=1}^{s}\alpha_{j}(B_{j}A_{j})x$. Crucially, upon entering a new stage $s$, we freeze both the backbone parameters $\theta_{B}$ and all previously introduced LoRA modules $\{\Delta\theta_{j}\}_{j=1}^{s-1}$. The optimization gradient is then applied only to the newly added module $\Delta\theta_{s}$ and the full set of scaling factors $\{\alpha_{j}\}_{j=1}^{s}$. This strategy offers a compelling dual advantage. First, it is computationally efficient by adaptively allocating resources according to task difficulty; new parameters are only introduced as needed for unsolved tasks, avoiding redundant computation on those already mastered. Second, the dynamic injection of new LoRA modules provides targeted plasticity for more challenging tasks without introducing negative transfer. By isolating updates within new modules while preserving the stable knowledge in the frozen backbone, our approach ensures that learning to solve difficult tasks does not disrupt previously acquired capabilities. In summary, DPS enables an intelligent allocation of model capacity, achieving a superior trade-off between efficiency and final performance. Pseudocode and implementation details are provided in Appendix A.3.
@@ 4 Experiments and Analysis
To comprehensively evaluate the effectiveness of our proposed ScaleZero model and the dynamic parameter scaling strategy, we designed a series of experiments. Section 4.1 aims to validate the performance of ScaleZero on several challenging heterogeneous multitask benchmarks (Atari 100k benchmark (Bellemare et al., 2013a), DeepMind Control Suite (Tunyasuvunakool et al., 2020), and Jericho(Hausknecht et al., 2020)), comparing it against specialized models trained independently for each task. Section 4.2 then focuses on quantitatively analyzing the specific gains in sample and computational efficiency afforded by our dynamic parameter scaling strategy (the combined method we called ScaleZero-DPS). All experiments are conducted in a purely online RL setting, without relying on any expert data, to test the model’s capabilities in a realistic learning scenario. Section 4.3 complements our benchmark results with a focused analysis of MoE, combining empirical evidence with theoretical justification to explain why MoE serves as an effective backbone for multitask learning. Implementation details are provided in Appendix A.

### v2

@@ 3.1 Diagnosing Plasticity Collapse in Multitask Planning
Figure 1: Plasticity collapse in the baseline (UniZero) on a multitask Atari benchmark. While simple tasks like Pong and Hero show stable learning, complex tasks such as Seaquest and ChopperCommand suffer a catastrophic performance collapse in later training (Top). This failure is precisely correlated with a sharp spike in the dormant neuron ratio of the transformer (Bottom Left) and an uncontrolled inflation of the latent state norm (Bottom Right), empirically validating the link between external performance and internal learning dynamics.
To empirically substantiate the challenges of representational bottlenecks and plasticity collapse, we conduct a quantitative analysis of a unified world model’s internal dynamics during multitask training. We track two key metrics indicative of network plasticity degradation: (1) $\textit{DormantRatio}(l)=\frac{1}{N_{l}}\sum_{i=1}^{N_{l}}\mathbf{1}!\left(|h_{i}^{l}|\leq\epsilon\right):$ This metric quantifies the proportion of neurons whose activation magnitude $|h_{i}^{l}|$ falls below a threshold $\epsilon$. A high ratio signifies a loss of active pathways and reduced network plasticity. (2) Latent State Norm: We compute the average L2 norm of the latent state ($|z_{t}|_{2}$). An uncontrolled inflation of this norm is a well-known indicator of training instability (Team et al., 2025).
Diagnostic experiments on UniZero, conducted across a multitask Atari8 suite (Appendix B), reveal a critical failure mode in complex learning scenarios. While simpler tasks like Pong achieve stable convergence, more complex ones such as Seaquest exhibit initial learning followed by a catastrophic performance collapse (Figure 1). This collapse coincides with a sharp increase in the dormant neuron ratio and an inflated latent state norm. The representational nature of this failure is corroborated by a marked decline in the feature effective rank (Dohare et al., 2024), indicating a degradation of the learned feature space (Figure 5). We attribute this failure mode to two intertwined factors: (1) Gradient Competition, where gradients from simpler tasks overwhelm those from complex ones, and (2) Representation Interference, where the shared network fails to maintain diverse features, leading to a representational bottleneck. Based on this diagnosis, our method tackles plasticity collapse from two complementary perspectives: (1) an internal, architectural solution, ScaleZero (Section 3.2), designed to mitigate interference within a single learning iteration. (2) an external, procedural strategy, Dynamic Parameter Scaling (Section 3.3), which optimizes resource allocation across the entire learning process.
3.2 Architecture Design of ScaleZero via Principled Exploration
@@
Figure 2: (a) A systematic exploration of the UniZero design space across five axes: task conditioning, encoder architecture, latent normalization, backbone design, and optimization. This investigation informs the design of our proposed ScaleZero model. (b) A conceptual diagram of Dynamic Parameter Scaling (DPS). DPS progressively expands model capacity by injecting LoRA adapters in stages, triggered by learning progress. This creates a curriculum of model, directing resources toward unsolved tasks while preserving existing knowledge.
@@
Figure 3: Performance impact of architectural modifications on the Atari8 multitask benchmark. This ablation across the UniZero design space reveals that replacing the Transformer backbone with a Mixture-of-Experts architecture yields the most significant and consistent performance gains. In contrast, other interventions, with the partial exception of SimNorm, provide marginal or inconsistent benefits. These results underscore the centrality of the MoE’s conditional computation in overcoming the limitations of a shared, dense backbone.
Motivated by the above diagnosis of plasticity collapse, we conduct a principled exploration of the UniZero design space to forge a robust solution on Atari8 multitask benchmark (see Appendix B for experimental details). We systematically evaluate five key design axes (Figure 2(a)) drawn from three core dimensions (input representation, model architecture, and optimization strategy): task conditioning, encoder architecture, latent normalization, backbone design, and optimization strategy. Specific implementation details are provided in Appendix A.3. The comprehensive results, summarized in Figure 3, reveal that the integration of a MoE backbone provides a uniquely significant and robust performance uplift, directly counteracting the failure modes identified in our initial diagnosis. Below, we detail the findings of each axis.
Explicit Task Conditioning. A shared model processing raw observations may face ambiguity in discerning the underlying task, especially when initial states are similar across environments (Rakelly et al., 2019; Sodhani et al., 2021). To test if providing an explicit task identity $k$ could resolve ambiguity, we concatenate a learnable task embedding $e_{k}$ to the latent state $z$. While this approach accelerates initial convergence on simpler tasks, it offers marginal long-term benefits for complex ones and fails to mitigate plasticity collapse, as evidenced by minimal improvement in metrics like the dormant neuron ratio (Figure 6). This indicates that such input-level task conditioning are insufficient to resolve deep-seated representational conflicts in online multitask RL.
Encoder Architecture (ResNet vs. ViT). We contrast a standard ResNet-like encoder in  Pu et al. (2024) with a Vision Transformer to assess the encoder’s impact on visual representation quality. On the Atari8 benchmark, the choice of encoder does not yield a significant performance difference. We attribute this to the limited visual heterogeneity in this specific task suite, but posit that the scalability and powerful feature extraction of ViT are critical for broader, more diverse multitask domains.
Latent Normalization (LayerNorm vs. SimNorm). To directly combat the observed latent norm inflation, we evaluate SimNorm (Hansen et al., 2023) against the standard LayerNorm (Ba et al., 2016). SimNorm, which projects latents onto a simplex to enforce a norm constraint, effectively stabilized training and prevented performance collapse, consistent with its reported success in less heterogeneous settings (Hansen et al., 2023). However, in our highly diverse multitask setup, this hard constraint appears to curtail representational expressiveness, leading to suboptimal final performance. This highlights a critical trade-off between stability and capacity for online heterogeneous RL training.
Mixture-of-Experts Backbone. The most impactful modification is replacing the dense feed-forward networks with sparse MoE layers in UniZero’s transformer. This conditional computation paradigm directly addresses the core challenges of multitask planning by mitigating the intertwined issues of representational interference and gradient conflict, as it routes task-specific computations and their corresponding gradient updates through distinct, specialized expert sub-networks. As validated in Figures 3 and 6, the MoE backbone yields substantial performance gains. We attribute this success to its inherent ability to mitigate gradient conflict, a mechanism we analyze in depth in Section 4.3. There, we provide both empirical validation and a theoretical analysis demonstrating that the upper bound on gradient conflict for an MoE layer is strictly lower than that of its dense counterpart.
@@
Multitask Gradient Correction. We also explored a dynamic gradient re-weighting scheme inspired by MoCo (Fernando et al., 2023) to directly mitigate gradient interference between tasks. While this approach showed promise on some tasks, it introduced significant computational overhead by increasing per-step training time by approximately 40%, and demonstrated inconsistent efficacy across our heterogeneous task suite. Given this unfavorable performance-cost trade-off, we defer the investigation of more efficient gradient correction methods to future work.
The ScaleZero Architecture. The result of our systematic investigation is ScaleZero, a novel architecture designed as a highly efficient and scalable world model. It is defined by the integration of three core components: (1) a ViT-like Encoder for powerful and scalable feature extraction; (2) an MoE Backbone that uses sparse, conditional computation to increase model capacity while mitigating task interference; and (3) an explicit Layer Normalization applied to the encoded latent state to ensure a robust balance between training stability and representational expressiveness. This core is complemented by a modular design that decouples task-specific encoders and heads, enabling flexible handling of diverse input and output modalities. As validated in Section 4, ScaleZero provides an effective blueprint for building generalist world models. Key differences from UniZero are detailed in Appendix A.2.
3.3 Dynamic Parameter Scaling for Efficient Multitask Learning
To address the second challenge in multitask planning—static resource allocation—we propose Dynamic Parameter Scaling (DPS). This strategy, illustrated in Figure 2(b), abandons the one-size-fits-all paradigm by dynamically coupling model capacity with learning progress. DPS operates on two synergistic principles: (1) adaptive task curation, which focuses computational effort exclusively on unsolved tasks, and (2) progressive capacity expansion, which injects new parameters only when needed. This combination creates a "curriculum of model complexity," directing resources precisely where they are most required. DPS employs dynamic, staged training to manage tasks and parameters.
  • Task Curation: Let $\mathcal{T}=\{\tau_{i}\}_{i=1}^{N}$ be the full set of tasks. A task $\tau_{i}$ is deemed "solved" once its performance metric, $\text{Metric}(\tau_{i})$, surpasses a threshold $\varepsilon_{i}$. At any time $t$, we maintain an active set of unsolved tasks, $\mathcal{U}_{t}\subseteq\mathcal{T}$. To maximize efficiency, both data collection and gradient updates are performed only for tasks within $\mathcal{U}_{t}$. Once solved, a task is removed from this set, ceasing all computational overhead associated with it.
  • Staged Capacity Expansion: The training proceeds in $S+1$ stages. Stage 0 (Warm-up) trains the base model $\theta_{B}$ for an initial $T_{0}$ iterations. Triggered by learning progress, each subsequent Stage $s\geq 1$ (Expansion) injects a new, independent LoRA module, $\Delta\theta_{s}$. This module adapts a target weight matrix of dimensions $d\times k$ via a low-rank decomposition $\Delta\theta_{s}=B_{s}A_{s}$, where $B_{s}\in\mathbb{R}^{d\times r}$, $A_{s}\in\mathbb{R}^{r\times k}$, and the rank $r\ll\min(d,k)$. The module’s update is scaled by a learnable scalar $\alpha_{s}$.
@@ 3.3 Dynamic Parameter Scaling for Efficient Multitask Learning
Adaptive Stage Transition Triggers. The transition from stage s1s-1 to $s$ is initiated when either of two conditions is met, balancing progress against a fixed budget: (1) Progress-based Trigger: A new stage begins if the number of newly solved tasks in the current stage reaches a quota $Q_{s}$. (2) Budget-based Trigger: To prevent stagnation on particularly arduous tasks, a transition is forced if the iteration count in the current stage exceeds a pre-allocated limit, $\lceil(T_{\max}-T_{0})/S\rceil$.
Optimization and Parameter Isolation. In any given stage $s$, the effective computation for a weight matrix $W_{0}\in\theta_{B}$ is expressed as $W^{(s)}(x)=W_{0}x+\sum_{j=1}^{s}\alpha_{j}(B_{j}A_{j})x$. Crucially, upon entering stage $s$, all prior parameters—the backbone $\theta_{B}$ and adapters $\{\Delta\theta_{j}\}_{j=1}^{s-1}$—are frozen. Optimization is thereby focused on the newly added LoRA module $\Delta\theta_{s}$ and the full set of scaling factors $\{\alpha_{j}\}_{j=1}^{s}$.
This strategy provides two decisive advantages. First, it implements a resource-efficient training curriculum, as new parameters are only introduced when required by the remaining unsolved tasks. Second, by isolating new learning within dedicated adapters, DPS provides targeted plasticity for difficult tasks while preventing catastrophic forgetting or negative transfer to previously mastered skills, whose knowledge is preserved in the frozen backbone. DPS thus achieves a superior trade-off between computational efficiency and final performance. (See Appendix A.4 for pseudocode).
@@ 4 Experiments and Analysis
Our experiments are designed to rigorously validate the two central claims of this work. First, in Section 4.1, we demonstrate that our proposed ScaleZero effectively overcomes the representational bottlenecks and plasticity collapse issue. We test its ability to achieve state-of-the-art performance as a single, generalist agent across a suite of heterogeneous multi-task benchmarks spanning visual, state-based, and text-based domains (Atari (Bellemare et al., 2013), DMControl (Tunyasuvunakool et al., 2020), and Jericho (Hausknecht et al., 2020)). Second, in Section 4.2, we quantify the efficiency gains of our DPS strategy, showing it successfully addresses the problem of static resource allocation. Finally, in Section 4.3, we provide a dedicated empirical and theoretical analysis of the Mixture-of-Experts mechanism, substantiating its critical role as the architectural cornerstone of ScaleZero’s success. All experiments are conducted in a purely online RL setting, without reliance on expert data, to assess true learning capability. Implementation details are provided in Appendix A.